# Weak convergence of Bayes estimators under general loss functions

### v1

@@ Abstract
Keywords: Asymptotic normality, Bernstein–von Mises theorem, Wasserstein distance, $L^{2}$-differentiability, Stein discrepancy, Bregman divergence.
@@ 1 Introduction
In Bayesian inference (see e.g. Berger,, 1985; Bernardo and Smith,, 1994; Gelman et al.,, 2014), the posterior distribution plays a central role by representing the updated beliefs about unknown parameters after observing data. Through Bayes’ theorem, it synthesizes prior knowledge with the observed data via the likelihood function. This unifying formulation provides a coherent framework for estimation, prediction, and decision-making under uncertainty. A foundational result in the asymptotic theory of Bayesian procedures is the Bernstein–von Mises (BvM) theorem, which traces back to the seminal insights of Laplace, (1820), and was later rigorously formalized in finite-dimensional settings by Bernstein, (1917) and von Mises, (1931), see also Giné and Nickl, (2016) and Ghosal and van der Vaart, (2017) for a modern account and extensions to semi- and nonparametric models. The classical BvM theorem asserts that, in a parametric model, under regularity conditions, the posterior distribution converges in total variation distance to a normal distribution centered at an efficient estimator (e.g. the maximum likelihood estimator) as the sample size grows. This convergence not only bridges Bayesian and frequentist paradigms but also implies that any mapping of the posterior distribution that is continuous with respect to total variation (e.g., the posterior mean or median) converges to the same mapping applied to the limiting normal distribution. In particular, it implies that Bayesian credible sets can admit an asymptotic frequentist interpretation as confidence sets, see Kleijn and van der Vaart, (2012) for further discussion.
Bayes estimators, defined as minimizers of expected posterior loss, can be viewed as such mappings. Hence, the BvM theorem offers a powerful framework for analyzing their asymptotic properties. The study of the large-sample behavior of Bayes estimators has long been a central theme in mathematical statistics, giving rise to a substantial body of literature, see e.g. Chao, (1970), Ibragimov and Has minskiĭ, (1973), Lhéritier and Iasnogorodski, (2005), as well as textbooks by Ibragimov and Has minskiĭ, (1981, Theorems 10.2 and 10.8) and van der Vaart, (1998, Theorem 10.8). These results establish that, under weak conditions on the model, Bayes estimators are both asymptotically normal and asymptotically efficient. A common assumption in the existing theory is that the loss function defining the Bayes estimator is translation invariant. More precisely, in a parametric setup, the loss function $\ell:\Theta\times\Theta\to[0,\infty)$, for a parameter space $\Theta\subseteq\mathbb{R}^{d}$, is required to satisfy
@@
for some function $\ell_{0}:\mathbb{R}^{d}\to[0,\infty)$. This structural condition encompasses several commonly used loss functions (Wang et al.,, 2022), most prominently the squared Euclidean distance, which yields the posterior mean as the Bayes estimator.
@@
However, many loss functions of practical relevance in modern applications do not satisfy the translation invariance condition (1). Notable examples include total variation distance, Stein discrepancies, intrinsic losses, Bregman divergences and Wasserstein distances (as well as its surrogates such as the Sinkhorn divergence). The latter have been a primary motivation for this work, as it has received substantial attention across disciplines owing to its ability to incorporate geometric information and to provide meaningful comparisons between probability distributions. Applications span fields such as image retrieval (Rubner et al.,, 2000), computer vision and image processing (Kolouri et al.,, 2019), generative machine learning (Arjovsky et al.,, 2017), econometrics and risk management (Feng and Schlogl,, 2018), natural language processing (Colombo et al.,, 2021), and biology (Schiebinger et al.,, 2019; Tameling et al.,, 2021; Bunne et al.,, 2023). In Bayesian statistics, the Wasserstein distance has also gained prominence, with uses ranging from predictive estimation (Matsuda and Strawderman,, 2021; Backhoff-Veraguas et al.,, 2022) to approximate computation (Bernton et al., 2019a, ; Ambrogioni et al.,, 2018) and sampling algorithms (Minsker et al.,, 2014; Srivastava et al.,, 2015, 2018). However, the theoretical understanding of the resulting Bayes estimators remains rather limited. For instance, Matsuda and Strawderman, (2021) analyzed the Bayesian predictive density for location–scale families (with a one-dimensional scale parameter) under the squared $2$-Wasserstein loss, showing that it reduces to the plug-in estimator based on the posterior mean, thereby falling within the well-understood framework of translation-invariant losses. In a nonparametric setting, Backhoff-Veraguas et al., (2022) established consistency of Bayes estimators with respect to $p$-Wasserstein losses ($p\geq 1$) under exponential moment conditions. By contrast, we consider parametric models and study not only consistency but, more importantly, the asymptotic distribution of Bayes estimators under mild moment conditions for general losses, including Wasserstein and translation-invariant ones. As an example, we show (in Section 5.1) that the Bayes estimator under the squared 2-Wasserstein loss is consistent and asymptotic normal for the Pareto family $\{P_{\vartheta}=\mathrm{Pareto}(1,\vartheta)\mid\vartheta\in[a,\infty)\}$ with some $a>2$. In this example, the consistency result of Backhoff-Veraguas et al., (2022, Theorem 4.24) does not apply, since $\int_{1}^{\infty}e^{\lambda_{0}\left\lVert x-x_{0}\right\rVert^{2}}\mathop{}\!\mathrm{d}P_{3}(x)=\infty,$ for all $\lambda_{0}>0$ and $x_{0}\geq 1$, which violates their exponential moment assumption.
@@
Motivated by the increasing relevance of such non-standard loss functions, we aim to develop a general theoretical framework for analyzing the asymptotic properties of Bayes estimators under broad classes of losses, particularly those that are not translation invariant. Our main contributions are as follows:
@@
    We establish consistency and derive asymptotic distributions for Bayes estimators associated with general loss functions that exhibit locally polynomial behavior, under fairly weak regularity assumptions.
@@
    We demonstrate that Bayes estimators under the squared 2-Wasserstein loss and Stein discrepancies are asymptotically efficient and asymptotically normal, thereby extending classical results to non-standard losses.
@@
    As an auxiliary contribution, we provide sufficient conditions for the usual differentiability of the loss function induced by the 2-Wasserstein distance, which is of independent interest.
Organization of the paper. Section 2 lays the foundation for our results, including a review of the BvM theorem in Section 2.1, and novel consistency results for Bayes estimators under general loss functions in Section 2.2. In Section 3, we present our main results on the weak convergence of Bayes estimators for losses that behave locally as polynomials. This section also contains a convergence result for the posterior expectation of loss functions, uniformly over compact sets, with its interpretation discussed in Section A.1. Section 4 specializes to loss functions that are locally quadratic. In Section 5, we provide several examples, including popular loss functions both within and beyond the classical framework (1). Even when the conditions of our theoretical framework are not satisfied, one may still expect a distributional limit, though possibly non-Gaussian, see Section 5.3. Numerical illustrations confirming the theoretical distributional limits of Bayes estimators are presented in Section 6. All proofs, along with an overview of different types of differentiability, and sufficient conditions for the differentiability of Wasserstein-type losses, are deferred to the supplementary material.
@@ 2.1 Bernstein–von Mises theorem and its implications
$E_{P^{n}_{\vartheta_{0}}}\left[\phi_{n}({\boldsymbol{X}})\right]\to 0\qquad\text{and}\qquad\sup_{\vartheta\in\Theta:\left\lVert\vartheta-\vartheta_{0}\right\rVert\geq\varepsilon}E_{P_{\vartheta}^{n}}\left[1-\phi_{n}({\boldsymbol{X}})\right]\to 0.$
@@ 2.2 Bayes estimator and its consistency
    $\ell(t,\vartheta)=g(\|t-\vartheta\|)$, where $g:[0,\infty)\to[0,\infty)$ is a non-decreasing function satisfying $g(x)=0$ if and only if $x=0$.
@@ 3 Weak convergence of Bayes estimators for locally polynomial losses
    There exist $\delta>0$ and constants $c_{3},c_{4}>0$ and $0<c_{5}<2$ such that $\ell\in\text{EB}(c_{3},c_{4},c_{5},B_{\delta}(\vartheta_{0}))$, as in Definition 1.
@@
If $\hat{\theta}_{n}$ is a Bayes estimator with respect to $\ell$ and $\hat{\theta}_{n}\stackrel{{\scriptstyle P^{n}_{\vartheta_{0}}}}{{\to}}\vartheta_{0}$, then $n^{1/2}\left(\hat{\theta}_{n}-\vartheta_{0}\right)\stackrel{{\scriptstyle D}}{{\to}}_{P^{n}_{\vartheta_{0}}}G.$
@@
Condition (ii) ensures that integrals over the sets $\{h\in H_{n}:\|h\|>M_{n}\}$ are asymptotically negligible by Proposition 7. This can be weakened to
@@
Losses of the form $\ell(t,\vartheta)=\|t-\vartheta\|^{p}$ satisfy the assumptions of Theorem 2 (see Propositions 7 and 8 and Section 2.2). By van der Vaart, (1998, Lemma 8.5), in this case $G=Y$, implying asymptotic efficiency of Bayes estimators under such losses. Further examples are given below and in the next section.
@@
Assume that at $\vartheta_{0}$, we have $D^{j}(\vartheta_{0})[u]=0$ if $j<k$ and $D^{k}(\vartheta_{0})[u]>0$, for every $u\in\mathbb{R}^{d}$. Then, by continuity (see Section C.3 for details), there exists $\varepsilon>0$ such that for all $t,\vartheta\in B_{\varepsilon}(\vartheta_{0})$, the upper and lower bounds in Theorem 2 (i) hold with $p=k$, i.e.
@@
The next lemma provides sufficient conditions under which Theorem 2 (iii) holds. It extends Proposition 8 (in Appendix B) to sequences of loss functions.
@@
In Lemma 1, condition (ii) links $\ell_{n}$ and $\ell_{0}$, ensuring convergence of $Z_{n}$ to $Z$. Condition (iii) controls the growth of the loss functions, guarantees boundedness on compact sets, and allows the use of Proposition 7. In this paper, Lemma 1 is applied with
@@
for which $\ell_{0}(t,h)$ is typically the directional Hadamard derivative of $\ell$ at $(\vartheta_{0},\vartheta_{0})$ in direction $(t,h)$. In this case, the weak limit $G$ minimizes, almost surely, the expected slope for $p=1$, the expected curvature for $p=2$, and so on, of $\ell$ at $(\vartheta_{0},\vartheta_{0})$ in direction $(G,h)$, with $h\sim\mathcal{N}_{d}(Y,I^{-1}_{\vartheta_{0}})$ and $Y\sim\mathcal{N}_{d}(0,I^{-1}_{\vartheta_{0}})$ (see also Section A.1).
@@ 4 Asymptotic efficiency of Bayes estimators under locally quadratic losses
    There exist $\delta>0$, $c_{2}>0$, and $0<a<2$ such that for all $t\in B_{\delta}(\vartheta_{0})$ and $\vartheta\in\Theta$,
    $|\xi(t,\vartheta)|\lesssim\exp\big(c_{2}(\|t\|^{a}+\|\vartheta\|^{a})\big),\text{ and }\|A(\vartheta)\|_{\mathsf{F}}\lesssim\exp\big(c_{2}\|\vartheta\|^{a}\big).$
The growth bounds on $\xi$ and $A$ in Assumption 3 are mild and are satisfied in many settings. The key requirement is the expansion (3). If $t\mapsto\ell(t,\vartheta)$ is thrice continuously differentiable for every $\vartheta\in\Theta$, then (3) follows from the second-order Taylor expansion whenever $\Theta\subseteq\mathbb{R}^{d}$ is convex. In this case, Assumption 3 holds with $c_{1}=3$ and $A(\vartheta)=\frac{1}{2}\nabla^{2}_{t}\ell(t,\vartheta)\big|_{t=\vartheta}$, provided $A(\vartheta_{0})$ is positive definite.
@@
$\|t-\vartheta\|^{2}\lesssim\ell(t,\vartheta)\lesssim\|t-\vartheta\|^{2}.$ (4)
The upper bound follows from
For the lower bound, let $\lambda_{\min}(\vartheta)$ denote the smallest eigenvalue of $A(\vartheta)$. By continuity of $G$ and positive definiteness of $A(\vartheta_{0})$, there exists $\delta>0$ such that
and $\eta_{\delta}$ is nonincreasing in $\delta$, we can choose $\varepsilon>0$ and $\eta>0$ such that $\eta_{\delta}\geq\eta>0$ for all $\delta<\varepsilon$. If $\delta<\varepsilon$ and $t,\vartheta\in B_{\delta}(\vartheta_{0})$,
Given the local expansion (3), the sequence of loss functions
@@
i.e., $\hat{\theta}_{n}$ is asymptotically efficient.

### v2

@@ Abstract
Keywords: Asymptotic normality, Bernstein–von Mises theorem, Wasserstein distance, Sinkhorn divergence, Stein discrepancy, Bregman divergence.
@@ 1 Introduction
In Bayesian inference (see e.g. Berger,, 1985; Bernardo and Smith,, 1994; Gelman et al.,, 2014), the posterior distribution plays a central role by representing the updated beliefs about unknown parameters after observing data. Through Bayes’ theorem, it synthesizes prior knowledge with the observed data via the likelihood function. This unifying formulation provides a coherent framework for estimation, prediction, and decision-making under uncertainty. A foundational result in the asymptotic theory of Bayesian procedures is the Bernstein–von Mises (BvM) theorem, which traces back to the seminal insights of Laplace, (1820), and was later rigorously formalized in finite-dimensional settings by Bernstein, (1917) and von Mises, (1931), see also Giné and Nickl, (2016); Castillo and Rousseau, (2015); Ghosal and van der Vaart, (2017); Durante et al., (2024) for a modern account and extensions to semi- and nonparametric models and non-gaussian limit distributions. The classical BvM theorem asserts that, in a parametric model, under regularity conditions, the posterior distribution converges in total variation distance to a normal distribution centered at an efficient estimator (e.g. the maximum likelihood estimator) as the sample size grows. This convergence not only bridges Bayesian and frequentist paradigms but also implies that any mapping of the posterior distribution that is continuous with respect to total variation (e.g., the posterior mean or median) converges to the same mapping applied to the limiting normal distribution. In particular, it implies that Bayesian credible sets can admit an asymptotic frequentist interpretation as confidence sets, see Kleijn and van der Vaart, (2012) for further discussion.
Bayes estimators, defined as minimizers of expected posterior loss, can be viewed as such mappings. The BvM theorem thus offers a powerful framework for analyzing their asymptotic properties. The canonical example is the posterior mean, arising under squared Euclidean loss, whose asymptotic properties are classical (LeCam,, 1953; Chao,, 1970; Ibragimov and Has minskiĭ,, 1973). Extensions (cf. Bickel and Yahav,, 1969) beyond this quadratic setting typically assume that the loss $\ell$ is translation invariant,
@@
for some $\ell_{0}:\mathbb{R}^{d}\to[0,\infty)$. An exception is Bunke and Milhaud, (1998), where translation invariance is imposed only on the global upper and lower bounds of the loss $\ell$ rather than on $\ell$ itself. A further common restriction is that the loss $\ell$ depends on a rescaled difference of parameters,
$\ell(t,\vartheta)=\ell_{n}(t,\vartheta)=\ell_{0}(\sqrt{n}(t-\vartheta))$ (2)
as considered in Ibragimov and Has minskiĭ, (1981); Jeganathan, (1982); van der Vaart, (1998); Lhéritier and Iasnogorodski, (2005). This rescaled-loss formulation (2) largely simplifies the asymptotic analysis. In particular, under (2), the rescaled estimator $\sqrt{n}(\hat{\theta}_{n}-\vartheta_{0})$ can be viewed as the minimizer of a posterior risk functional whose integrand no longer depends on the sample size $n$. This invariance eliminates a major technical obstacle in establishing asymptotic results. However, except in the special case where the loss is positively homogeneous, the estimator defined through (2) does not coincide with the standard Bayes estimator associated with a fixed loss function.
@@
By contrast, many loss functions of practical relevance in modern applications neither satisfy the translation invariance condition (1) nor admit global upper and lower bounds that are themselves translation invariant. Notable examples include total variation distance, Stein discrepancies, intrinsic losses, Bregman divergences and Wasserstein-type distances (as well as their surrogates such as the Sinkhorn divergence). The latter have been a primary motivation for this work, as they recently receive substantial attention across disciplines owing to its ability to incorporate geometric information and to provide meaningful comparisons between probability distributions. Applications span fields such as image retrieval (Rubner et al.,, 2000), computer vision and image processing (Kolouri et al.,, 2019), generative machine learning (Arjovsky et al.,, 2017), econometrics and risk management (Feng and Schlogl,, 2018), natural language processing (Colombo et al.,, 2021), and biology (Schiebinger et al.,, 2019; Tameling et al.,, 2021; Bunne et al.,, 2023). In Bayesian statistics, the Wasserstein distance has also gained prominence, with uses ranging from predictive estimation (Matsuda and Strawderman,, 2021; Backhoff-Veraguas et al.,, 2022) to approximate computation (Bernton et al., 2019a, ; Ambrogioni et al.,, 2018) and sampling algorithms (Minsker et al.,, 2014; Srivastava et al.,, 2015, 2018). However, the theoretical understanding of the resulting Bayes estimators remains rather limited. For instance, Matsuda and Strawderman, (2021) analyzed the Bayesian predictive density for location–scale families (with a one-dimensional scale parameter) under the squared $2$-Wasserstein loss, showing that it reduces to the plug-in estimator based on the posterior mean, thereby falling within the well-understood framework of translation-invariant losses. In a nonparametric setting, Backhoff-Veraguas et al., (2022) established consistency of Bayes estimators with respect to $p$-Wasserstein losses ($p\geq 1$) under exponential moment conditions. By contrast, we consider parametric models and study not only consistency but, more importantly, the asymptotic distribution of Bayes estimators under mild moment conditions for general losses, including Wasserstein and translation-invariant ones. As an example, we show (in Section 5.1) that the Bayes estimator under the squared 2-Wasserstein loss is consistent and asymptotic normal for the Pareto family $\{P_{\vartheta}=\mathrm{Pareto}(1,\vartheta)\mid\vartheta\in[a,\infty)\}$ with some $a>2$. In this example, the consistency result of Backhoff-Veraguas et al., (2022, Theorem 4.24) does not apply, since $\int_{1}^{\infty}e^{c\left\lVert x-x_{0}\right\rVert^{2}}\mathop{}\!\mathrm{d}P_{3}(x)=\infty,$ for all $c>0$ and $x_{0}\geq 1$, which violates their exponential moment assumption.
@@
Motivated by the growing importance of such non-standard loss functions, we develop a general theoretical framework for analyzing the asymptotic behavior of Bayes estimators under broad classes of losses, in particular, those that are not translation invariant. For such losses, the structural simplifications afforded by (1) or (2) are no longer available, and the asymptotic analysis requires substantially more delicate arguments, even though, at a high level, the overall proof strategy remains conceptually aligned with the classical works of LeCam, (1953); Chao, (1970); Ibragimov and Has minskiĭ, (1973). Our main contributions are as follows:
@@
    We establish consistency (Proposition 2) and derive asymptotic distributions (Theorem 2) for Bayes estimators associated with general loss functions that exhibit locally polynomial behavior, under comparatively mild regularity conditions. A key structural insight is that the posterior risk, viewed as a stochastic process indexed by the local parameter (i.e. a suitable local rescaling of the parameter), admits a nondegenerate limit typically characterized by the Hadamard directional derivative of the loss function (Proposition 3). This derivative naturally emerges as the effective limiting loss, and the asymptotic distribution of the Bayes estimator is therefore governed by a location functional of a Gaussian limit experiment evaluated with respect to this derivative. In this way, the Hadamard directional derivative of the loss plays a central role in linking local posterior asymptotics with the limiting optimization problem.
@@
    We show that Bayes estimators under squared $2$-Wasserstein loss (Section 5.1) and under Stein discrepancies (Section 5.2) are asymptotically normal and achieve asymptotic efficiency, thereby extending classical optimality results beyond the translation-invariant framework.
@@
    As an auxiliary contribution, we provide sufficient conditions (Theorems 4 and 5 in Supplement A.3) for classical differentiability of loss functions induced by the $2$-Wasserstein distance. These are of independent interest for the theoretical analysis of Wasserstein-based procedures.
Organization of the paper. Section 2 lays the foundation for our results, including a review of the BvM theorem in Section 2.1, and novel consistency results for Bayes estimators under general loss functions in Section 2.2. In Section 3, we present our main results on the weak convergence of Bayes estimators for losses that behave locally as polynomials. This section also contains a convergence result for the posterior expectation of loss functions, uniformly over compact sets, with its interpretation discussed in Supplement A.1. Section 4 specializes to loss functions that are locally quadratic. In Section 5, we provide several examples, including popular loss functions both within and beyond the classical framework (1). Even when the conditions of our theoretical framework are not satisfied, one may still expect a distributional limit, though possibly non-Gaussian, see Section 5.3. Numerical illustrations confirming the theoretical distributional limits of Bayes estimators are presented in Section 6. A discussion of potential extensions of the proposed framework is given in Section 7. All proofs, along with an overview of different types of differentiability, and sufficient conditions for the differentiability of Wasserstein-type losses, are deferred to the supplementary material.
@@ 2.1 Bernstein–von Mises theorem and its implications
$\mathbb{E}_{P^{n}_{\vartheta_{0}}}\left[\phi_{n}({\boldsymbol{X}})\right]\to 0\qquad\text{and}\qquad\sup_{\vartheta\in\Theta:\left\lVert\vartheta-\vartheta_{0}\right\rVert\geq\varepsilon}\mathbb{E}_{P_{\vartheta}^{n}}\left[1-\phi_{n}({\boldsymbol{X}})\right]\to 0.$
@@ 2.2 Bayes estimator and its consistency
    $\ell(t,\vartheta){\,\geq\,}g(\|t-\vartheta\|)$, where $g:[0,\infty)\to[0,\infty)$ is a non-decreasing function satisfying $g(x)=0$ if and only if $x=0$.
@@ 3 Weak convergence of Bayes estimators for locally polynomial losses
    There exist $\delta>0$ and constants $c_{3},c_{4}>0$ and $0<c_{5}<2$ such that $\ell\in\text{EB}(c_{3},c_{4},c_{5},B_{\delta}(\vartheta_{0}))$, see Definition 1.
@@
If $\hat{\theta}_{n}$ is a Bayes estimator with respect to $\ell$ and $\hat{\theta}_{n}\stackrel{{\scriptstyle P^{n}_{\vartheta_{0}}}}{{\to}}\vartheta_{0}$, then $n^{1/2}\bigl(\hat{\theta}_{n}-\vartheta_{0}\bigr)\stackrel{{\scriptstyle D}}{{\to}}_{P^{n}_{\vartheta_{0}}}G.$
@@
Condition (ii) ensures that integrals over the sets $\{h\in H_{n}:\|h\|>M_{n}\}$ are asymptotically negligible by Proposition 7 (in Supplement B). This can be weakened to
@@
Losses of the form $\ell(t,\vartheta)=\|t-\vartheta\|^{p}$ satisfy the assumptions of Theorem 2 (see Section 2.2, and Propositions 7 and 8 in Supplement B). By van der Vaart, (1998, Lemma 8.5), in this case $G=Y$, implying asymptotic efficiency of Bayes estimators under such losses. Further examples are given below and in the next section.
@@
Assume that at $\vartheta_{0}$, we have $D^{j}(\vartheta_{0})[u]=0$ if $j<k$ and $D^{k}(\vartheta_{0})[u]>0$, for every $u\in\mathbb{R}^{d}$. Then, by continuity (see Supplement C.3 for details), there exists $\varepsilon>0$ such that for all $t,\vartheta\in B_{\varepsilon}(\vartheta_{0})$, the upper and lower bounds in Theorem 2 (i) hold with $p=k$, i.e.
@@
The next lemma provides sufficient conditions under which Theorem 2 (iii) holds. It extends Proposition 8 (in Supplement B) to sequences of loss functions.
@@
In Lemma 1, condition (ii) links $\ell_{n}$ and $\ell_{0}$, ensuring convergence of $Z_{n}$ to $Z$. Condition (iii) controls the growth of the loss functions, guarantees boundedness on compact sets, and allows the use of Proposition 7 (in Supplement B). In this paper, Lemma 1 is applied with
@@
for which $\ell_{0}(t,h)$ is typically the directional Hadamard derivative of $\ell$ at $(\vartheta_{0},\vartheta_{0})$ in direction $(t,h)$. In this case, the weak limit $G$ minimizes, almost surely, the expected slope for $p=1$, the expected curvature for $p=2$, and so on, of $\ell$ at $(\vartheta_{0},\vartheta_{0})$ in direction $(G,h)$, with $h\sim\mathcal{N}_{d}(Y,I^{-1}_{\vartheta_{0}})$ and $Y\sim\mathcal{N}_{d}(0,I^{-1}_{\vartheta_{0}})$ (see also Supplement A.1).
@@ 4 Asymptotic efficiency of Bayes estimators under locally quadratic losses
    There exist $\delta,c_{2},c_{3}>0$ and $0<c_{4}<2$ such that $\ell\in\text{EB}(c_{2},c_{3},c_{4},B_{\delta}(\vartheta_{0}))$, see Definition 1.
The growth bounds in Assumption 3 (i)(ii) are rather mild and hold in many settings. The key requirement is the expansion (4). If $t\mapsto\ell(t,\vartheta)$ is thrice continuously differentiable for every $\vartheta\in\Theta$, then (4) follows from the second-order Taylor expansion whenever $\Theta\subseteq\mathbb{R}^{d}$ is convex. In this case, Assumption 3 holds with $c_{1}=3$ and $A(\vartheta)=\frac{1}{2}\nabla^{2}_{t}\ell(t,\vartheta)\big|_{t=\vartheta}$, provided $A(\vartheta_{0})$ is positive definite.
@@
$\|t-\vartheta\|^{2}\lesssim\ell(t,\vartheta)\lesssim\|t-\vartheta\|^{2},$ (5)
see Section C.4.1. Given the local expansion (4), the sequence of loss functions
@@
i.e. $\hat{\theta}_{n}$ is asymptotically efficient.